# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ 4 Quasi-continuum modes
The huge potential of quasi-mode SARIs to govern instability in accretion disks becomes clear when plotting composite spectra for several wavenumbers, where the Doppler range moves for every $m$. In Fig. 3, we show spectra for modes with wavenumbers $m\in\{-3,-2,\ldots,3\}$ and wavenumbers $k/|m|=50$ for the SARIs and $k=50$ for the MRI, recreating Fig. 20 of Goedbloed & Keppens (2022) using Legolas. We use the combined efforts of the QR algorithm to find the quasi-continuum edges and the Arnoldi shift-invert algorithm to fill the 2D regions. The colour code adopted here will be used in the remainder of this paper: the $m=0$ MRI in black and the $m<0$ counter- and $m>0$ co-rotating SARIs in red and blue, respectively, to give the Doppler shift its usual connotation. We find an excellent correspondence with the $W_{\text{com}}\leq 10^{-12}$ regions delimited by ROC. The Alfvén continua (red and black) and Doppler range (gray) are indicated for each value of $m$, and it is clear that each quasi-continuum is limited to the Doppler range. The furthest edges of the $m=\pm 3$ continua fall out of this frame at around $\sigma=-3.5$ and $3.8$, respectively. With Legolas, we additionally obtain the discrete SARI modes clustering towards the continua, as well as stable Alfvénic modes (visible on the real axis for $m=\pm 1$). Note again the up-down symmetry of the spectrum, a feature of ideal MHD (Goedbloed et al., 2019). This figure underlines that a typical equilibrium has a huge range of unstable perturbations over many wavenumbers, with perturbations that can be radially localised anywhere in the disk! If a linear MHD view suffices to understand the cause of turbulence, this abundance of entire 2D regions of near-eigenmode SARIs that grow exponentially and experience differental rotation in the disk is a much more convincing argument than the existence of a few isolated unstable MRIs.
@@ 5 SARI and MRI in accretion disk equilibria
Before exploring the instabilities further numerically, it is useful to state the MRI/SARI instability criterion and analytical maximum growth rate predictions derived by Goedbloed & Keppens (2022). This upper bound for growth, obtained from the global problem (their Eq. (84)), improve on an upper bound derived under the WKB approximation (Balbus & Hawley, 1992):
@@
Here, the squared epicyclic frequency $\kappa_{e}^{2}=4\Omega^{2}+r\left(\Omega^{2}\right)^{\prime}=\Omega^{2}$ for our Keplerian rotation profile. Note that the Alfvén frequency $\omega_{A}=\left(\frac{m}{r}B_{\theta}+kB_{z}\right)/\sqrt{\rho}$ contains both wavenumbers $k$ and $m$, and it can now go through zero and become negative for $m<0$, in contrast to the situation for the MRI. If $\omega_{A}$ vanishes and the lower bound is violated, the modes should stabilise. The same should happen if the upper bound is violated because the field gets too strong or the wavenumbers too large – physically, magnetic tension then stabilises modes with wavelengths that are sufficiently small. Equation (9) can also be differentiated with respect to $\omega_{A}^{2}$ to obtain the maximum growth rate over all wavenumers at a fixed radius,
@@
The dimensionless parameters $\epsilon$, $\beta$, $\mu_{1}$ and $\delta$ defined in Sect. 2.1 completely determine an ideal MHD equilibrium with cylindrical flow and gravity, representing an accretion disk. Goedbloed & Keppens (2022) take $v_{A1}=B_{1}=\epsilon\ll 1$ and $p_{1}\sim\epsilon^{2}\beta\ll 1$, and $k^{2}r^{2}\gg m^{2}$ so that the equilibrium state represents a near-Keplerian ($\Omega_{1}\approx 1$), weakly magnetised, thin accretion disk (vertical scale height $H\sim c_{S}=\frac{\gamma}{2}v_{A1}^{2}\beta\ll 1$ (Pringle, 1981)). Super-Alfvénic rotation then follows, since $v_{A1}\ll 1\sim m\Omega_{1}$. These assumptions allow for an analytical treatment of the global eigenvalue problem, but a numerical code like Legolas is not limited to those regimes. Nevertheless, we will find that in most cases, the analytical growth rate predictions are still sufficient to understand stability of the modes (Sect. 5.4).
@@
  • Case (a): $\epsilon=0.0141$, $\mu_{1}=1$, $\beta=100$, $\delta=1,k=50$, corresponding to a weak field and pitch angle $\varphi=45^{\circ}$ (see Sect. 2.1);
  • Case (b): $\epsilon=0.045$, $\mu_{1}=10$, $\beta=10$, $\delta=1,k=50$, corresponding to a slightly stronger, predominantly toroidal field ($\varphi\approx 85^{\circ}$),
where one or more parameters are varied. Both sets of parameters have a sizeable $B_{\theta}$ component – indeed, there is a growing set of evidence that significant toroidal fields may be generated in accretion disks (Das et al., 2018; Begelman et al., 2022). The main differences between the two cases are then equipartition between poloidal/toroidal field versus dominant poloidal field, and weak field to slightly less weak field. In both cases, $\Omega_{1}\approx 0.987$.
Legolas works in dimensionless units, which are fixed in all runs by the following default choice: $B_{u}=10$ G for the magnetic field, $L_{u}=10^{9}$ cm as length unit, and $T_{u}=10^{6}$ K for temperature. The spectra of Figs. 1, 2, 3 and 11 are given in these dimensionless values. The growth rates in Figs. 48 are additionally normalised by the inner disk rotation frequency $\Omega_{1}$. All spectra in those Figures were calculated at a resolution of $200$ base gridpoints.
@@ 5.1 Vertical localisation
Figure 4 shows how the growth rates vary with $k$ for Case (b) ($\varphi=85^{\circ}$), which is more illustrative here than Case (a). The growth rates of the $m=0$ MRI are shown in black, the $m>0$ and $m<0$ SARIs in blue and red, respectively, with the inner SARIs in solid lines and the outer branch dotted. One immediate observation is that all instabilities have a clear cut-off wavelength and reach maximum growth for intermediate wavelengths. Both the cut-off wavenumber $k_{\text{cut-off}}$ and wavenumber of maximal growth $k_{\text{max}}$ are ordered by $m$ and significantly increased for counter-rotating $m<0$ SARIs, which allow for much smaller unstable vertical wavelengths, as was also found by Begelman et al. (2022). Maximum growth rates are comparable for the MRI and inner SARI, but the outer SARIs consistently have lower growth rates. The latter growth rate curves largely follow those of the MRI and inner SARIs, but scaled down with a factor of two horizontally and a factor of around three vertically. The reason for this scaling will be explained in 5.4. Strikingly, for $m<0$ SARIs, there is an intermediate value of $k$ for which the growth rates vanish before the maximum is reached. This is most prominent for $m=-10$, where the inner branch is stabilised at $k=100$ and the outer branch at $k=50$. At these wavenumbers, the Alfvén frequency vanishes locally, i.e. $\omega_{A}(r_{*})=0$, at the corotation location $r_{*}$, which violates the SARI instability criterion given by Eq. (10). This highlights that the magnetic field is essential for instability, and instability does not prevail for those wavenumbers where the wave is directed perpendicularly to the field ($\mathbf{k}\cdot\mathbf{B}=0$). It is important to bear in mind that the range of allowed $k$ is limited in a realistic, 2D disk with a vertical extent and hence a quantisation on $k$ or at least an upper limit for vertical wavelengths. Hence, the range of $k$ where the modes are physical is determined by $H\sim c_{S}$, which gives in this case $k\gtrsim 30$. Case (a) (not shown) has much smaller values of $k_{\text{max}}$ and $k_{\text{cut-off}}$ because of the stronger vertical field component. The variation with $m$ is similar to Case (b) but less pronounced: because the field has a larger vertical component ($B_{z1}=B_{\theta 1}$), the influence of the azimuthal wavenumber $m$ is reduced and one needs to go to more negative $m$ to significantly increase $k_{\text{cut-off}}$.
Figure 4: Normalised growth rates for MRI (black, dashdot) and SARI (inner: solid, outer: dashed) over a range of vertical wavenumbers $k$ in Case (b) (dominant azimuthal field). Both $k_{\text{cut-off}}$ and $k_{\text{max}}$ increase as $m$ decreases. Outer growth rates are scaled-down versions of the inner curves. Cusps in the growth rate curves, e.g. at $k=100$ for the inner $m=-10$ SARI, signify local stabilisation if $\omega_{A}$ vanishes.
@@ 5.2 Field orientation
Figure 5 (a) shows the growth rates for Case (a) (weak field and $\beta=100$). The MRI and low-$m$ SARI are the fastest-growing modes whenever there is a significant $B_{z}$-component present, with inner and outer SARI growth rates decreasing with $|m|$. All growth rates drop significantly as the field becomes more toroidal, but the SARI growth rates remain finite whereas the MRI is stabilised completely as $\varphi\rightarrow 90^{\circ}$: with $B_{z}\rightarrow 0$ and $m=0$ there is no way for the instability to couple to the toroidal component, and $\omega_{A}=0$ everywhere (Balbus & Hawley, 1992). For a purely toroidal field, the growth rates increase with $|m|$. This is understood from Fig. 4: an equivalent figure for fixed $k$ and varying $m$ would show an equivalent growth rate curve for $m$ with an $m_{\text{max}}$ and $m_{\text{cut-off}}$. Hence, for values of $|m|$ below $m_{\text{max}}$, the growth rates increase with $|m|$. As in Fig. 4, for every $m<0$ there is a value of $\varphi$ where $\omega_{A}(r_{*})=0$ near the inner or outer boundary and as a result the respective inner or outer mode that resonates with this frequency is locally stabilised.
Figure 5 (b) of Case (b) (slightly stronger field and $\beta=10$) shows quite a different situation: at $k=50$, both MRI and SARI are suppressed when the vertical field component dominates, and they are only activated for a sizeable toroidal component. The ‘cut-off’ inverse pitch angle is seen to decrease as $m$ becomes increasingly negative: the relative strength of the poloidal field allowed for instability increases. Comparing with Case (a), we conclude that for $\epsilon=0.014$, the vertical component is able to support at $k=50$ instability, but $\epsilon=0.045$ is too strong unless $m$ is very negative. The cut-off being less sharp for $m=\pm 10$ SARIs is a resolution effect. For a purely toroidal field, the MRI is again suppressed and SARIs reach comparably large growth rates, again increasing with $|m|$ in this range of azimuthal wavenumbers. Again, $m<0$ SARIs can be stabilised locally. The maximum growth rates of all instabilities are comparable to those of Case (a), but the stronger azimuthal field increases the SARI growth rates in the limit $\varphi\rightarrow 90^{\circ}$.
Figure 5: Normalised growth rates for MRI (black, dashdot) and SARI (inner: solid, outer: dashed) for purely vertical ($\varphi=0^{\circ}$) to purely azimuthal ($\varphi=90^{\circ}$) background field for Case (a) and Case (b) at $k=50$. Note how a slight increase in field strength in Panel (b) greatly reduces the allowed poloidal component. MRI stabilises in a purely toroidal field, SARIs become more unstable.
@@ 5.3 Magnetisation
Two of the essential assumptions for SARI modes (and for the standard MRI in Balbus & Hawley (1991)) are a weak field, i.e. $v_{A1}\ll 1$ and thin disk with scale height $H\sim c_{S1}=\frac{\gamma}{2}v_{A1}^{2}\beta\ll 1$, so that near-Keplerian disks rotate super-Alfvénically and the background is close to incompressible. We now investigate fields from weak to dynamically relevant, modeling disks with near-Keplerian to strongly sub-Keplerian rotation speeds.
Figure 7 shows MRI and SARI growth rates as a function of the Alfvén speed $v_{A1}=\epsilon$ normalised by the inner rotation speed $v_{\theta 1}=r_{1}\Omega_{1}$, with the most unstable inner and outer SARIs again denoted by solid and dashed lines, respectively. Panel (a) and (b) again show Case (a) with $\varphi=45^{\circ}$ and Case (b) with a dominant toroidal field, respectively, both with $k=50$ and $\beta>10$. An immediate result is that for every $k,m$ there is a bounded range of field strengths at which instability operates. For Case (a), all growth rates are maximal for $v_{A1}\approx 0.02v_{\theta 1}$, well in the extremely super-Alfvénic regime, but for Case (b) the range of $v_{A1}$ in which instability prevails for the different $m$ is much larger, with $v_{A1}/v_{\theta 1}$ close to unity for $m<0$. The $m=-3$ SARI almost doubles the range of instability compared to the MRI at this value of $k$. In both panels, the value of $v_{A1\text{max}}$ and $v_{A1\text{cut-off}}$ shift with the value of $m$ in a distinctive way. In Panel (a), both $v_{A1\text{max}}$ and $v_{A1\text{cut-off}}$ decrease as the wavenumbers go from $m<0$ to $m>0$. In Panel (b), the shifts are more pronounced and do not obey the same ordering: remarkably, the $m=-10$ SARI is cut off at significantly weaker fields than other $m<0$ SARIs, and its $v_{A1\text{max}}$ corresponds roughly with that of the MRI. This pattern will be explained from the SARI instability criterion in Sect. 5.4. Note that in both cases the outer SARIs again behave very similarly to the inner SARIs, but with smaller growth rates and stricter cut-off. In Case (b), the outer $m=-10$ SARI is stabilised, as $k=50$ is the vertical wavenumber where $\omega_{A}(r_{2})=0$ (this is also clear from Fig. 4). Note that the normalisation factor $\Omega_{1}$ is itself dependent on $\epsilon$ (Sect. 2.1). This parameter can become much smaller than its Keplerian value $\Omega_{1}=1$ for the strongest fields in Case (b), going down to $\Omega_{1}<0.3$. For such sub-Keplerian disks, the inner disk is magnetically supported instead of rotationally.
@@
The field strength can also be considered relative to the gas pressure through varying the plasma-beta. Although the standard MRI was first obtained for disks threaded by a weak, subthermal field (Balbus & Hawley, 1991), is has been shown to prevail in suprathermal disks as well (Das et al., 2018), motivated by simulations where a strong suprathermal toroidal field is generated ($\beta\sim 0.1$ to $1$) (Bai & Stone, 2013). Figure 8 (a) shows how the MRI and SARI growth rates depend on plasma-beta for our Case (a) with equal poloidal and toroidal field components. The MRI is the most unstable mode, closely followed by the inner SARI, with growth rates decreasing with $|m|$. ±mplus-or-$\pm m$ SARI growth rates almost coincide for $\beta<1$ but are separated for subthermal fields, which becomes more apparent as $|m|$ increases. For these equilibria, suprathermal magnetic fields ($\beta>1$) produce slightly larger growth rates than subthermal fields ($\beta<1$). The outer SARIs show a similar trend, but ±mplus-or-$\pm m$ growth rates almost coincide for subthermal fields. They diverge, however, in the limit $\beta\rightarrow 0$, where the co-rotating modes retain a finite growth rate but the counter-rotating modes have vanishing growth rate. Note that the $m=-10$ outer mode is stabilised completely for $\beta\rightarrow 0$, while the $m=10$ outer mode is briefly stabilised and then reappears. However, a closer inspection of the spectra for these equilibria (not shown) reveals that these modes are not SARIs: the slow continua are separated from the Alfvén continua and have their proper overlap, and the outer eigenfunctions are not confined to the region between the resonances with the Alfvén continuum. They could be related to the hybrid Alfvén/slow modes described in Das et al. (2018). Figure 8 (b) shows quite a different behaviour for Case (b) with a slightly stronger dominant toroidal field. All modes show a large variation in growth rates and the MRI is not the most unstable mode over most of the $\beta$ range. All growth rates become small for $\beta\rightarrow 0$. Note the sudden drop in growth of the $m=10$ modes at large values of $\beta$, while other growth rates drastically increase. This is a result of the normalisation, where $\Omega_{1}$ depends critically on $\beta$. When the disk becomes strongly sub-Keplerian, $\Omega_{1}\ll 1$. It is expected that for values of $\beta$ where $\Omega_{1}\rightarrow 0$, the growth rates will remain finite and suddenly vanish like the $m=10$ modes. In contrast to most of our results shown above, growth rates of the outer $m=-3$ SARI are slightly larger than those of the inner SARI over almost all of the parameter range. Indeed, the wavenumber $k=50$ is close to that where the $m=-3$ inner mode is locally stabilised (Fig. 4), and hence it has a reduced growth rate. The $m=-10$ outer mode is again fully stabilised at this wavenumber.

### v2

@@ 4 Quasi-continuum modes
The huge potential of quasi-mode SARIs to govern instability in accretion disks becomes clear when plotting composite spectra for several wavenumbers, where the Doppler range moves for every $m$. In Fig. 3, we show spectra for modes with wavenumbers $m\in\{-3,-2,\ldots,3\}$ and wavenumbers $k/|m|=50$ for the SARIs and $k=50$ for the MRI, recreating Fig. 20 of Goedbloed & Keppens (2022) using Legolas. We use the combined efforts of the QR algorithm to find the quasi-continuum edges and the Arnoldi shift-invert algorithm to fill the 2D regions. The colour code adopted here will be used in the remainder of this paper: the $m=0$ MRI in black and the $m<0$ counter- and $m>0$ co-rotating SARIs in red and blue, respectively, to give the Doppler shift its usual connotation. We find an excellent correspondence with the $W_{\text{com}}\leq 10^{-12}$ regions delimited by ROC. The Alfvén continua (red and black) and Doppler range (gray) are indicated for each value of $m$, and it is clear that each quasi-continuum is limited to the Doppler range. The furthest edges of the $m=\pm 3$ continua fall out of this frame at around $\sigma=-3.5$ and $3.8$, respectively. With Legolas, we additionally obtain the discrete SARI modes clustering towards the continua, as well as stable Alfvénic modes (visible on the real axis for $m=\pm 1$). Note again the up-down symmetry of the spectrum, a feature of ideal MHD (Goedbloed et al., 2019). This figure underlines that a typical equilibrium has a huge range of unstable perturbations over many wavenumbers, with perturbations that can be radially localised anywhere in the disk! If a linear MHD view suffices to explain the cause of turbulence, this abundance of entire 2D regions of near-eigenmode SARIs that grow exponentially and experience differential rotation in the disk is a much more convincing argument than the existence of a few isolated unstable MRIs.
@@ 5 SARI and MRI in accretion disk equilibria
Before exploring the instabilities further numerically, it is useful to state the MRI/SARI instability criterion and analytical maximum growth rate predictions derived by Goedbloed & Keppens (2022). This upper bound for growth, obtained from the global problem (their Eq. (84)), improves on an upper bound derived under the WKB approximation (Balbus & Hawley, 1992):
@@
Here, the squared epicyclic frequency $\kappa_{e}^{2}=4\Omega^{2}+r\left(\Omega^{2}\right)^{\prime}=\Omega^{2}$ for our Keplerian rotation profile. Note that the Alfvén frequency $\omega_{A}=\left(\frac{m}{r}B_{\theta}+kB_{z}\right)/\sqrt{\rho}$ contains both wavenumbers $k$ and $m$, and it can now go through zero and become negative for $m<0$, in contrast to the situation for the MRI. If $\omega_{A}$ vanishes and the lower bound is violated, the modes should stabilise. The same should happen if the upper bound is violated because the field gets too strong or the wavenumbers too large – physically, magnetic tension then stabilises modes with wavelengths that are sufficiently small. Equation (9) can also be differentiated with respect to $\omega_{A}^{2}$ to obtain the maximum growth rate over all wavenumbers at a fixed radius:
@@
The dimensionless parameters $\epsilon$, $\beta$, $\mu_{1}$ and $\delta$ defined in Sect. 2.1 completely determine an ideal MHD equilibrium with cylindrical flow and gravity, representing an accretion disk. Goedbloed & Keppens (2022) take $v_{A1}=B_{1}=\epsilon\ll 1$ and $p_{1}\sim\epsilon^{2}\beta\ll 1$, and $k^{2}r^{2}\gg m^{2}$ so that the equilibrium state represents a near-Keplerian ($\Omega_{1}\approx 1$), weakly magnetised, thin accretion disk (vertical scale height $H\sim c_{S}=\frac{\gamma}{2}v_{A1}^{2}\beta\ll 1$ (Pringle, 1981)). Super-Alfvénic rotation then follows, since $v_{A1}\ll 1\sim m\Omega_{1}$. These assumptions allow for an analytical treatment of the global eigenvalue problem, but a numerical code like Legolas is not limited to those regimes. Nevertheless, we will find that in most cases, the analytical growth rate predictions are still sufficient to understand the stability of the modes (Sect. 5.4).
@@
  • Case (a): $\epsilon=0.0141$, $\mu_{1}=1$, $\beta=100$, $\delta=1$, and $k=50$, corresponding to a weak field and pitch angle $\varphi=45^{\circ}$ (see Sect. 2.1); or
  • Case (b): $\epsilon=0.045$, $\mu_{1}=10$, $\beta=10$, $\delta=1$, and $k=50$, corresponding to a slightly stronger, predominantly toroidal field ($\varphi\approx 85^{\circ}$),
where one or more parameters are varied. Both sets of parameters have a sizeable $B_{\theta}$ component – indeed, there is a growing set of evidence that significant toroidal fields may be generated in accretion disks (Das et al., 2018; Begelman et al., 2022). The main differences between the two cases are then an equipartition between the poloidal and toroidal fields versus dominant poloidal field, and a weak field versus a slightly less weak field. In both cases, $\Omega_{1}\approx 0.987$.
Legolas works in dimensionless units, which are fixed in all runs by the following default choices: $B_{u}=10$ G for the magnetic field, $L_{u}=10^{9}$ cm as the length unit, and $T_{u}=10^{6}$ K for temperature. The spectra of Figs. 1, 2, 3 and 11 are given in these dimensionless values. The growth rates in Figs. 48 are additionally normalised by the inner disk rotation frequency $\Omega_{1}$. All spectra in those Figures are calculated at a resolution of $200$ base gridpoints.
@@ 5.1 Vertical localisation
Figure 4 shows how the growth rates vary with $k$ for Case (b) ($\varphi=85^{\circ}$), which is more illustrative here than Case (a). The growth rates of the $m=0$ MRI are shown in black, and those of the $m>0$ and $m<0$ SARIs in blue and red, respectively, with the inner SARIs in solid lines and the outer branch dotted. One immediate observation is that all instabilities have a clear cut-off wavelength and reach maximum growth for intermediate wavelengths. Both the cut-off wavenumber $k_{\text{cut-off}}$ and the wavenumber of maximal growth $k_{\text{max}}$ are ordered by $m$ and significantly increased for counter-rotating $m<0$ SARIs, which allow for much smaller unstable vertical wavelengths, as was also found by Begelman et al. (2022). Maximum growth rates are comparable for the MRI and inner SARI, but the outer SARIs consistently have lower growth rates. The latter growth rate curves largely follow those of the MRI and inner SARIs, but scaled down by a factor of 2 horizontally and a factor of around 3 vertically. The reason for this scaling will be explained in 5.4. Strikingly, for $m<0$ SARIs, there is an intermediate value of $k$ for which the growth rates vanish before the maximum is reached. This is most prominent for $m=-10$, where the inner branch is stabilised at $k=100$ and the outer branch at $k=50$. At these wavenumbers, the Alfvén frequency vanishes locally, i.e. $\omega_{A}(r_{*})=0$, at the corotation location $r_{*}$, which violates the SARI instability criterion given by Eq. (10). This highlights that the magnetic field is essential for instability, and instability does not prevail for those wavenumbers where the wave is directed perpendicularly to the field ($\mathbf{k}\cdot\mathbf{B}=0$). It is important to bear in mind that the range of allowed $k$ is limited in a realistic, 2D disk with a vertical extent and hence a quantisation on $k$ or at least an upper limit for vertical wavelengths. Hence, the range of $k$ where the modes are physical is determined by $H\sim c_{S}$, which gives in this case $k\gtrsim 30$. Case (a) (not shown) has much smaller values of $k_{\text{max}}$ and $k_{\text{cut-off}}$ because of the stronger vertical field component. The variation with $m$ is similar to Case (b) but less pronounced: because the field has a larger vertical component ($B_{z1}=B_{\theta 1}$), the influence of the azimuthal wavenumber $m$ is reduced and one needs to go to more negative $m$ to significantly increase $k_{\text{cut-off}}$.
Figure 4: Normalised growth rates for MRI (black, dashdot) and SARI (inner: solid, outer: dashed) over a range of vertical wavenumbers $k$ in Case (b) (dominant azimuthal field). Both $k_{\text{cut-off}}$ and $k_{\text{max}}$ increase as $m$ decreases. The outer growth rates are scaled-down versions of the inner curves. Cusps in the growth rate curves, e.g. at $k=100$ for the inner $m=-10$ SARI, signify local stabilisation if $\omega_{A}$ vanishes.
@@ 5.2 Field orientation
Figure 5 (a) shows the growth rates for Case (a) (weak field and $\beta=100$). The MRI and low-$m$ SARI are the fastest-growing modes whenever there is a significant $B_{z}$-component present, with inner and outer SARI growth rates decreasing with $|m|$. All growth rates drop significantly as the field becomes more toroidal, but the SARI growth rates remain finite whereas the MRI is stabilised completely as $\varphi\rightarrow 90^{\circ}$: with $B_{z}\rightarrow 0$ and $m=0$ there is no way for the instability to couple to the toroidal component, and $\omega_{A}=0$ everywhere (Balbus & Hawley, 1992). For a purely toroidal field, the growth rates increase with $|m|$. This is understood from Fig. 4: an equivalent figure for fixed $k$ and varying $m$ would show an equivalent growth rate curve for $m$ with an $m_{\text{max}}$ and $m_{\text{cut-off}}$. Hence, for values of $|m|$ below $m_{\text{max}}$, the growth rates increase with $|m|$. As in Fig. 4, for every $m<0$ there is a value of $\varphi$ at which $\omega_{A}(r_{*})=0$ near the inner or outer boundary and as a result the respective inner or outer mode that resonates with this frequency is locally stabilised.
Figure 5 (b) of Case (b) (slightly stronger field and $\beta=10$) shows quite a different situation: at $k=50$, both MRI and SARI are suppressed when the vertical field component dominates, and they are only activated for a sizeable toroidal component. The ‘cut-off’ inverse pitch angle is seen to decrease as $m$ becomes increasingly negative: the relative strength of the poloidal field allowed for instability increases. Comparing with Case (a), we conclude that for $\epsilon=0.014$, the vertical component is able to support instability at $k=50$ instability, but $\epsilon=0.045$ is too strong unless $m$ is very negative. The cut-off being less sharp for $m=\pm 10$ SARIs is a resolution effect. For a purely toroidal field, the MRI is again suppressed and SARIs reach comparably large growth rates, again increasing with $|m|$ in this range of azimuthal wavenumbers. Again, $m<0$ SARIs can be stabilised locally. The maximum growth rates of all instabilities are comparable to those of Case (a), but the stronger azimuthal field increases the SARI growth rates in the limit $\varphi\rightarrow 90^{\circ}$.
Figure 5: Normalised growth rates for MRI (black, dashdot) and SARI (inner: solid, outer: dashed) for purely vertical ($\varphi=0^{\circ}$) to purely azimuthal ($\varphi=90^{\circ}$) background field for Case (a) and Case (b) at $k=50$. Note how a slight increase in field strength in Panel (b) greatly reduces the allowed poloidal component. MRI stabilises in a purely toroidal field; SARIs become more unstable.
@@ 5.3 Magnetisation
Two of the essential assumptions for SARI modes (and for the standard MRI in Balbus & Hawley (1991)) are a weak field, i.e. $v_{A1}\ll 1$ and a thin disk with scale height $H\sim c_{S1}=\frac{\gamma}{2}v_{A1}^{2}\beta\ll 1$, so that near-Keplerian disks rotate super-Alfvénically and the background is close to incompressible. We now investigate fields from weak to dynamically relevant, modeling disks with near-Keplerian to strongly sub-Keplerian rotation speeds.
Figure 7 shows the MRI and SARI growth rates as a function of the Alfvén speed $v_{A1}=\epsilon$ normalised by the inner rotation speed $v_{\theta 1}=r_{1}\Omega_{1}$, with the most unstable inner and outer SARIs again denoted by solid and dashed lines, respectively. Panel (a) and (b) again show Case (a) with $\varphi=45^{\circ}$ and Case (b) with a dominant toroidal field, respectively, both with $k=50$ and $\beta>10$. An immediate result is that for every $k,m$ there is a bounded range of field strengths at which instability operates. For Case (a), all growth rates are maximal for $v_{A1}\approx 0.02v_{\theta 1}$, well in the extremely super-Alfvénic regime, but for Case (b) the range of $v_{A1}$ in which instability prevails for the different $m$ is much larger, with $v_{A1}/v_{\theta 1}$ close to unity for $m<0$. The $m=-3$ SARI almost doubles the range of instability as compared to the MRI at this value of $k$. In both panels, the values of $v_{A1\text{max}}$ and $v_{A1\text{cut-off}}$ shift with the value of $m$ in a distinctive way. In Panel (a), both $v_{A1\text{max}}$ and $v_{A1\text{cut-off}}$ decrease as the wavenumbers go from $m<0$ to $m>0$. In Panel (b), the shifts are more pronounced and do not obey the same ordering: remarkably, the $m=-10$ SARI is cut off at significantly weaker fields than other $m<0$ SARIs, and its $v_{A1\text{max}}$ corresponds roughly with that of the MRI. This pattern will be explained from the SARI instability criterion in Sect. 5.4. Note that in both cases the outer SARIs again behave very similarly to the inner SARIs, but with lower growth rates and stricter cut-off. In Case (b), the outer $m=-10$ SARI is stabilised, as $k=50$ is the vertical wavenumber where $\omega_{A}(r_{2})=0$ (this is also clear from Fig. 4). Note that the normalisation factor $\Omega_{1}$ is itself dependent on $\epsilon$ (Sect. 2.1). This parameter can become much smaller than its Keplerian value $\Omega_{1}=1$ for the strongest fields in Case (b), going down to $\Omega_{1}<0.3$. For such sub-Keplerian disks, the inner disk is magnetically supported instead of rotationally.
@@
The field strength can also be considered relative to the gas pressure through varying the plasma-beta. Although the standard MRI was first obtained for disks threaded by a weak, subthermal field (Balbus & Hawley, 1991), it has been shown to prevail in suprathermal disks as well (Das et al., 2018), motivated by simulations where a strong suprathermal toroidal field is generated ($\beta\sim 0.1$ to $1$) (Bai & Stone, 2013). Figure 8 (a) shows how the MRI and SARI growth rates depend on the plasma-beta for our Case (a) with equal poloidal and toroidal field components. The MRI is the most unstable mode, closely followed by the inner SARI, with growth rates decreasing with $|m|$. ±mplus-or-$\pm m$ SARI growth rates almost coincide for $\beta<1$ but are separated for subthermal fields, which becomes more apparent as $|m|$ increases. For these equilibria, suprathermal magnetic fields ($\beta>1$) produce slightly higher growth rates than subthermal fields ($\beta<1$). The outer SARIs show a similar trend, but the ±mplus-or-$\pm m$ growth rates almost coincide for subthermal fields. They diverge, however, in the limit $\beta\rightarrow 0$, where the co-rotating modes retain a finite growth rate but the counter-rotating modes have vanishing growth rate. Note that the $m=-10$ outer mode is stabilised completely for $\beta\rightarrow 0$, while the $m=10$ outer mode is briefly stabilised and then reappears. However, a closer inspection of the spectra for these equilibria (not shown) reveals that these modes are not SARIs: the slow continua are separated from the Alfvén continua and have their proper overlap, and the outer eigenfunctions are not confined to the region between the resonances with the Alfvén continuum. They could be related to the hybrid Alfvén/slow modes described in Das et al. (2018). Figure 8 (b) shows quite a different behaviour for Case (b) with a slightly stronger dominant toroidal field. All modes show a large variation in growth rates and the MRI is not the most unstable mode over most of the $\beta$ range. All growth rates become low for $\beta\rightarrow 0$. Note the sudden drop in growth of the $m=10$ modes at large values of $\beta$, while other growth rates drastically increase. This is a result of the normalisation, where $\Omega_{1}$ depends critically on $\beta$. When the disk becomes strongly sub-Keplerian, $\Omega_{1}\ll 1$. It is expected that for values of $\beta$ where $\Omega_{1}\rightarrow 0$, the growth rates will remain finite and suddenly vanish like those for the $m=10$ modes. In contrast to most of our results shown above, growth rates of the outer $m=-3$ SARI are slightly higher than those of the inner SARI over almost the entire parameter range. Indeed, the wavenumber $k=50$ is close to that where the $m=-3$ inner mode is locally stabilised (Fig. 4), and hence it has a reduced growth rate. The $m=-10$ outer mode is again fully stabilised at this wavenumber.